How many milliliters of rubbing alcohol (isopropyl alcohol) weighs 500. grams? The density of isopropyl alcohol is 0.785 g/mL.


1
Expert's answer
2021-05-17T04:08:34-0400

We know that,


"Density = \dfrac{Mass}{Volume}"

"Volume = \dfrac{Mass}{Density}"

"Volume = \dfrac{500}{0.785}"

"Volume = 636.94mL"